concept checkpoints The bolted connection is subjected to a tensile force of P 62 kN. The diameter of the bolt is d = 28 mm. Determine the average shear stress in the bolt (in MPa). enter Enter your answer (without units) in the text box 01-201
Answer #1

average shear stress \sigma =\frac{P}{2A}

\sigma =\frac{62000}{2\times (\frac{\pi }{4}\times 28^{2})}

\sigma =50.344 Mpa